OpenStudy (anonymous):

Rewrite with only sin x and cos x. sin 3x

OpenStudy (anonymous):

@hartnn

OpenStudy (anonymous):

@iPwnBunnies

OpenStudy (anonymous):

medal!!

OpenStudy (anonymous):

i just did it and got 2 sin x cos2x + cos x

OpenStudy (anonymous):

but the cosx might be wrong

OpenStudy (anonymous):

it might be sin^3x..?

OpenStudy (tkhunny):

Really? \(\sin(3x) = \sin(x + 2x) = \sin(x)\cos(2x)+\cos(x)\sin(2x)\) Now what?

OpenStudy (anonymous):

\[\sin 3x=\sin \left( 2x+x \right)=\sin 2x \cos x+\cos 2x \sin x\]

OpenStudy (anonymous):

i dont get it

OpenStudy (anonymous):

i dont get what i have to do next

OpenStudy (tkhunny):

\(\sin(2x) = ???\) \(\cos(2x) = ???\) There is a reasona why you studied all those identities? Now is the time to prove it.

OpenStudy (anonymous):

\[\sin 2x=\sin \left( x+x \right)=\sin x \cos x+\cos x \sin x=2\sin x \cos x\] \[\cos 2x=\cos \left( x+x \right)=\cos x~\cos x+\sin x~\sin x=\cos ^2x-\sin ^2x\]

OpenStudy (anonymous):

substitute and simplify

OpenStudy (anonymous):

thank you @surjithayer

OpenStudy (anonymous):

now i add them?

OpenStudy (anonymous):

what do i do now?.... @surjithayer

OpenStudy (tkhunny):

?? What are you not seeing? Did you go through the steps so far provided? We have seen none of YOUR work. This would be a good time to demonstrate your algebra skills. Apply the identities. Simplify where you can. There is nothing magic about it.

OpenStudy (anonymous):

@mathslover

OpenStudy (anonymous):

I DO NOT SEE WHAT I HAVE TO DO NEXT. WHAT DONT YOU UNDERSTAND?

OpenStudy (anonymous):

@tkhunny

OpenStudy (anonymous):

@uri

OpenStudy (tkhunny):

Did you substitute the identities? If not, that is what you do next. Did you simplify where possible? If not, that is what you do next. Did you complete the exercise? If not, that is what you do next. Did you move on to the next problem? If not, that is what you do next - unless you want to take a lunch break, or something.

OpenStudy (anonymous):

i got it right, soooo bye

OpenStudy (tkhunny):

Woo-hoo!!! Too bad you showed us NONE of your work. Very, VERY hard to help you if you show us nothing. I am glad you got it.